# CodeSSM: Towards State Space Models for Code Understanding

### v2

@@ Abstract
Although transformers are widely used for various code-specific tasks, they have some significant limitations. In this paper, we investigate State Space Models (SSMs) as a potential alternative to transformers for code understanding tasks, such as code retrieval, classification, and clone detection. Previous research has already demonstrated that SSMs are more compute-efficient than transformers. In our work, we show that SSMs are also more sample-efficient and can effectively extrapolate to longer contexts (beyond the pretraining context) during fine-tuning. Through comprehensive experiments, we demonstrate that SSMs could serve as a viable alternative to transformers for code understanding tasks, while addressing some of the major limitations associated with transformers.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) have been known to perform well on various code intelligence tasks. One of the main reasons for this improvement is the pretraining - finetuning paradigm used to train transformer models, which is possible due to the ease of parallelization of attention-based transformer architectures. Under this paradigm, the transformer model is initially trained on a large corpus of unlabeled data using a self-supervised training objective. This is followed by a supervised finetuning process on a smaller labeled dataset.
However, the performance gain offered by the transformer models comes with some trade-offs, such as quadratic complexity, a substantial data requirement, and high inference costs. While methods such as linear attention (Katharopoulos et al., 2020) and sparse attention (Guo et al., 2022b; Zaheer et al., 2020; Condevaux and Harispe, 2022) have been proposed to address computational inefficiency, the practical gains from these methods are limited (Yang et al., 2025; Qin et al., 2022).
@@
One-dimensional Convolutional Neural Networks (1D CNNs) are a fast and positional embedding-free alternative to transformers. 1D CNNs, when trained using the pretraining-finetuning approach, show improved performance on some tasks but fail to capture long-range dependencies in input data (Tay et al., 2021).
@@
State-space models (SSMs)Gu et al. (2021, 2022b, 2022a) represent an alternative that addresses the limitations of transformers and 1D CNNs. SSMs leverage the strengths of 1D CNNs while effectively capturing long-range dependencies with linear-time complexity. Additionally, SSMs are position-aware, eliminating the need for positional embedding.
@@
In this work, we explore the application of SSMs to the domain of code understanding. To achieve this, we pretrain an encoder-only model based on SSM, which we call CodeSSM, on a small code dataset with masked language modeling (MLM) (Devlin et al., 2019). We evaluated our model on multiple code understanding benchmarks. We also investigate various training aspects, such as the effects of positional embedding, dropout, bidirectional SSMs, and varying pretraining context lengths. We compare the CodeSSM with two transformer models that we train – (1) BertCoder: based on Bert architecture (Devlin et al., 2019) and (2) RoCoder: based on RoFormer architecture Su et al. (2024). We assess the performance of our model against several high-performing transformers as well.
@@
Our study demonstrates that SSMs exhibit strong sample efficiency in code language modeling tasks. Notably, CodeSSM achieves over 50% accuracy in masked language modeling (MLM) with fewer than 3,000 training steps.
@@
Additionally, the code language modeling capabilities of CodeSSM also transfer well to downstream tasks as CodeSSM outperforms transformers on multiple code understanding tasks when pretrained under similar conditions. CodeSSM is also competitive with (and even surpasses) transformer models trained on significantly larger datasets and with more complex pretraining objectives. Furthermore, another advantage of SSMs is their ability to extrapolate to much larger contexts than those encountered during pretraining, effectively addressing a key limitation of the transformer architecture.
@@
Our objective in training the CodeSSM model is not to achieve state-of-the-art performance on coding benchmarks, but to explore and document the strengths and limitations of SSMs in understanding code. This goal also motivates our usage of a small dataset and MLM pretraining which allows us to quickly experiment with different model configurations and compare them under similar training settings. Although we observe improvements in performance with larger pretraining data, we will leave the exploration of very large training datasets and more complex training objectives for future work.
@@
  • We conduct the first systematic evaluation of State Space Models (SSMs) for code understanding, introducing a specialized model, which we term CodeSSM. Our analysis reveals that CodeSSM achieves robust performance across multiple benchmarks, consistently outperforming attention-based models while requiring substantially less pretraining data.
  • Our study presents the first empirical evidence that State Space Models (SSMs) achieve superior sample efficiency in code-understanding tasks compared to transformers. CodeSSM substantially outperforms transformer baselines in data-constrained scenarios with only a few thousand training samples.
@@
  • We observe that including even a small dropout mask in the CodeSSM layer degrades the performance of the model even on very small benchmarks. CodeSSM also performs very well on the adversarial text-code search task. Together these observations indicate that SSMs exhibit robust semantic understanding capabilities.
@@ 2.1 Architecture
CodeSSM-dropout. Dropout layers (Srivastava et al., 2014) prevent over-fitting in neural networks. So, to investigate over-fitting in CodeSSM, we create a variant with two dropout masks, positioned after each multiplication operations (see Figure 1). We refer to this variant as CodeSSM-do.
CodeF and CodeSSM-Uni. Recent research Lee-Thorp et al. (2022) questioned the necessity of self-attention in transformers. To investigate the significance of SSM, we conducted a similar ablation study by replacing SSM blocks in CodeSSM with DFT (see Figure 1), creating a variant we call CodeF. Furthermore, we analyze the importance of bidirectional processing via CodeSSM-Uni (Unidirectional Gated SSM), a variant of CodeSSM that removes flip operations from the CodeSSM layer.
@@ 2.2 Pretraining
Dataset. We initially trained CodeSSM and a BERT-like transformer model, which we refer to as BertCoder, on 1.8M text-code pairs from the CodeSearchNet (CSN) dataset. The small pretraining dataset allows us to quickly experiment with various model configurations. Both models were trained with a large context window of 2048, enabling them to be fine-tuned on benchmarks with lengths shorter than 2048. We refer to the CodeSSM model trained with the CSN dataset as CodeSSM-base. Additionally, we also train CodeSSM variants such as CodeSSM-pos, CodeSSM-do, CodeF and CodeSSM-Uni (see 2.1) on the same dataset and context length.
Context Length and Length Extrapolation. To investigate the length extrapolation capability of SSMs, we pretrain the CodeSSM model with a 256-length context window. Since we cannot change the context window for BertCoder once pretrained with a smaller context, we pretrained a RoFormer model Su et al. (2024), which is similar to the BERT model but with Rotatory Positional Embeddings (RoPE), with a 256-length context window. We refer to this model as RoCoder.
CSN dataset separates docstring and code, which creates an issue when they are concatenated and subsequently truncated. This can result in one modality (text) being much longer than the other modality (code). We also observed this discrepancy results in reduction in performance in tasks involving two sequences (refer to column 3 of Table 2).
To alleviate this issue, we use a subset of 1.8M samples from the StarCoder dataset (Li et al., 2023) to train the models with 256-length context (referred to as CodeSSM-sc-256). This subset contains 300k samples from the six programming languages (PLs) in the CSN dataset  [^2] Go, Java, Javascript, PHP, Python, Ruby. We also train CodeSSM on this dataset with a context length of 2048 (referred to as CodeSSM-sc-2048).
Data Scaling. To study the impact of large datasets on the downstream performance of CodeSSM, we also train it on a large subset of the StarCoder dataset with 4.8M samples (referred as CodeSSM-sc-L). This subset contains 600k samples from 8 PLs (C and C# along with CSN PLs).
@@ 3 Experiments
In this section, we describe the benchmarks on which we evaluate CodeSSM and the baselines with which we compare our model.
@@ 3.1 Baselines
We trained two transformer-based models as baselines: a large context (2048) BERT model to compare SSMs’ large context abilities and a small context RoFormer model to compare the length extrapolation abilities of transformers and SSMs. The models being compared are trained under the same settings: same dataset, same objective, context length, and batch size.
Additionally, we also mention the performance of some well-performing transformer models in each task to place CodeSSM in the context of SOTA encoder-only transformer models. However, it should be noted that these transformer models are trained on a much larger dataset and with a mixture of multiple training objectives (see Table 1).
@@ 3.2 Benchmarks
Retrieval tasks. There are two retrieval tasks on which we evaluate our models, NLCodeSearch (AdvTest) Lu et al. (2021) and Stackoverflow Question Answer (SQA). In NLCodeSearch, a description in natural language is provided, and the goal is to find the source code that corresponds to that description. To evaluate a model’s ability to generalize, function names and variable names in the test sets are replaced with generic tokens, for example, the function name with func. This task is difficult because the model has to locate the correct code throughout the test set. In SQA, a StackOverflow question is given and the goal is to retrieve the highest upvoted answer. This dataset is derived from the original StackOverflow dataset  [^3] https://www.kaggle.com/datasets/stackoverflow/stacksample/data. The average length of question and answer are 1.2k and 1.4k respectively.
Sequence classification tasks. Sequence classification includes two vulnerability detection tasks, Devign Zhou et al. (2019) and DiverseVul Chen et al. (2023a), as well as a complexity prediction task Jeon et al. (2023). Vulnerability detection is a binary classification task where the goal is to identify whether a given code contains security vulnerabilities. The Devign benchmark Zhou et al. (2019) is a balanced dataset, while DiverseVul consists of only around 5% vulnerable samples. So, we use accuracy for the former and F1 macro for the latter. Complexity prediction Jeon et al. (2023) is a small benchmark with only 3613 training samples to predict the algorithmic complexity of a given Java code among 7 different labels.
Token classification task. In the type-inference Jesse and Devanbu (2022) task, the model has to predict the type (such as integer, boolean, etc.) of each token in a given code. The label consists of 50K most frequent (built-in and user-defined) types, while the remaining types have been replaced with the UNK type. During the evaluation, predicting the UNK token is considered incorrect.
@@
For all tasks, we maintain the same context length as specified in the original benchmarks. All models are fine-tuned using a context length of 512 for text-to-code search, clone detection, type inference, and DiverseVul. For the Devign benchmark, the models are trained with a context length of 400. In SQA and complexity prediction, we use a context length of 1024.
@@
Table 3: Results on SQA and APPS benchmark in terms of MRR. The best performance is in bold and other noteworthy results are underlined.
@@ 4 Results
We present the performance of our model on various benchmarks described in Section 3.2
@@ 4.1 Retrieval Tasks
NLCodeSearch(AdvTest). Column 3 of Table 2 presents results of various models evaluated on NLCodeSearch benchmark. Among the models trained on the CSN dataset, CodeSSM-base (with a context size of 2048) outperforms other models, including CodeBERT, which was pretrained with significantly more data. CodeSSM-base with a context length of 256 performs poorly compared to other variants – we explain the reason for this in Section 2.2. CodeSSM-sc-256 shows remarkable performance and length extrapolation to double the length on which it was trained and is competitive with RoCoder and CodeSSM-sc-2048.
@@
StackOverflowQA. Table 3 presents the performance of CodeSSM in conjunction with the baseline models trained by us, BertCoder and RoCoder, on the SQA benchmark. SQA allows us to evaluate the models with longer contexts. CodeSSM-sc with a pretraining context length of 256 achieves strong results – it successfully extrapolates to contexts four times longer than those seen during pretraining. Furthermore, CodeSSM-sc outperforms BertCoder (pretrained with a context length of 2048) by 6 points, despite being pretrained with a context length eight times shorter. CodeSSM also surpasses RoCoder, highlighting its superior ability to extrapolate to longer contexts and capture long-range dependencies. Furthermore, we observe that increasing the pretraining context length leads to further improvements in CodeSSM’s performance.
@@ 4.2 Sequence classification tasks
Devign. Column 3 of Table 4 shows the results of the various models on the devign benchmark. CodeSSM outperforms BertCoder and RoCoder as well as CodeBERT, GraphCodeBERT, and CodeT5. Interestingly, BertCoder, pretrained with a longer context, also outperforms CodeBERT, showcasing the importance of a longer context for transformers. However, in the case of CodeSSM, the smaller context model performs much better. Notably, RoCoder performs poorly, potentially due to over-fitting, as the devign dataset is very small (only 21k training samples). RoPE, used in RoCoder, is also prone to positional overfitting Liu et al. (2024b); Goel et al. (2025).
DiverseVul. Column 2 of Table 4 shows that CodeSSM and its variants outperform BertCoder and RoCoder on the DiverseVul benchmark, which is a larger vulnerability detection dataset compared to Devign. It shows that CodeSSM can perform better than transformers even on a larger dataset. CodeSSM even outperforms CodeBERT, GraphCodeBERT, and CodeT5, which are trained on a much larger dataset and have multiple pretraining objectives (see Table 1).
Complexity Prediction. Table 5 shows that CodeSSM-sc-256 outperforms all the transformer models even the ones trained with longer context length (BertCoder) and larger data and more pretraining objectives (CodeBERT, GraphCodeBERT, UnixCoder). Additionally, CodeSSM-sc-2048 shows very strong performance as pretraining on longer context length improves the performance further on this task.

### v3

@@ Abstract
Although transformers dominate many code-specific tasks, they have significant limitations. This paper explores State Space Models (SSMs) as a promising alternative for code understanding tasks such as retrieval, classification, and clone detection. We introduce CodeSSM, the first SSM-based model trained on code corpora to assess its effectiveness. Our results demonstrate that SSMs are more sample-efficient and can extrapolate to longer contexts beyond the pretraining length. Extensive experiments show that SSMs offer a viable alternative to transformers, addressing several their limitations. Additionally, CodeSSM reduces memory usage by up to 64% compared to transformers at a context length of 2048, with greater savings as context length grows.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) have been known to perform well in various applications. One of the main reasons for this improvement is the pretraining - fine-tuning paradigm used to train transformer models, which is possible due to the ease of parallelization of attention-based transformer architectures. Under this paradigm, the transformer model is initially trained on a large corpus of unlabeled data using a self-supervised training objective, followed by a supervised fine-tuning process on a smaller labeled dataset.
However, the performance gain offered by the transformer models comes with some trade-offs, such as quadratic complexity, a substantial data requirement, and high inference costs. Although methods such as linear attention (Katharopoulos et al., 2020) and sparse attention (Guo et al., 2022b; Zaheer et al., 2020; Condevaux and Harispe, 2022) have been proposed to address computational inefficiency, the practical gains from these methods are limited (Yang et al., 2025; Qin et al., 2022).
@@
One-dimensional Convolutional Neural Networks (1D CNNs) are a fast and positional embedding-free alternative to transformers. 1D CNNs, when trained using the pretraining-finetuning approach, show improved performance on some tasks, but do not capture long-range dependencies in input data (Tay et al., 2021).
Figure 1: Memory consumption and throughput of CodeSSM and RoCoder model during inference.
@@
State-space models (SSMs)Gu et al. (2021, 2022b, 2022a) represent an alternative that addresses the limitations of transformers and 1D CNNs. SSMs leverage the strengths of 1D CNNs while effectively capturing long-range dependencies with linear-time complexity. In addition, SSMs are position-aware, eliminating the need for positional embedding.
@@
In this work, we explore the application of SSMs to the domain of code understanding. To achieve this, we pretrain an encoder-only model based on SSM, which we call CodeSSM, on a small code dataset with masked language modeling (MLM) (Devlin et al., 2019). We evaluated our model on multiple code understanding benchmarks. We also investigate various training aspects, such as the effects of positional embedding, dropout, bidirectional SSMs, and varying pretraining context lengths. We compare CodeSSM with two transformer models that we train – (1) BertCoder: based on the Bert architecture (Devlin et al., 2019) and (2) RoCoder: based on the RoFormer architecture Su et al. (2024). Moreover, we compare our model with Zamba, a hybrid SSM transformer model based on the Zamba2 architecture Glorioso et al. (2024). We also assess the performance of our model against several high-performing transformers, which are trained with larger data and complex pretraining objectives.
@@
Our study demonstrates that SSMs exhibit strong sample efficiency in code modeling tasks. In particular, CodeSSM achieves more than 50% precision in Masked Language Modeling (MLM) with fewer than 3,000 training steps. Furthermore, we show that CodeSSM consumes up to 64% less memory compared to transformers and is significantly faster (see Figure˜1). The efficiency in terms of speed and memory consumption in CodeSSM (compared to transformers) improves with increasing input length.
@@
Additionally, the code language modeling capabilities of CodeSSM also transfer well to downstream tasks, as CodeSSM outperforms transformers on multiple code understanding tasks when pretrained under similar conditions. CodeSSM is also competitive with (and even surpasses) transformer models trained on significantly larger datasets and with more complex pretraining objectives. Furthermore, another advantage of SSMs is their ability to extrapolate to much larger contexts than those encountered during pretraining, effectively addressing a key limitation of the transformer architecture.
@@
Our goal in training the CodeSSM model is not to achieve state-of-the-art results on coding benchmarks, but to explore and document the strengths and limitations of SSMs for code understanding. This motivated our use of a small dataset and MLM pretraining, enabling rapid experimentation with and comparison of different model configurations under consistent training conditions. Although we observed that larger pretraining datasets improved performance, the investigation of extensive datasets and more complex training objectives is left for future work.
@@
  • We present the first systematic evaluation of State-Space Models (SSMs) for code understanding, introducing a specialized model, which we term CodeSSM. Our analysis reveals that CodeSSM achieves robust performance on retrieval, classification and clone detection tasks, consistently outperforming attention-based models while requiring substantially less pretraining data and memory.
  • Our study presents the first empirical evidence that SSMs achieve superior sample efficiency in code-understanding tasks compared to transformers. CodeSSM substantially outperforms transformer baselines in data-constrained scenarios with only a few thousand training samples.
@@
  • We observe that including even a small dropout mask in the CodeSSM layer degrades the performance of the model even on very small benchmarks. CodeSSM also performs very well on the adversarial text-code search task. Together, these observations indicate that SSMs exhibit robust semantic understanding capabilities.
The code is available here.
@@ 2.1 Architecture
CodeSSM-dropout. Dropout layers (Srivastava et al., 2014) prevent overfitting in neural networks. So, to investigate overfitting in CodeSSM, we create a variant with two dropout masks, positioned after each multiplication operation (see Figure˜2). We refer to this variant as CodeSSM-do.
CodeF and CodeSSM-Uni. Recent research questioned the necessity of self-attention in transformers Lee-Thorp et al. (2022). To investigate the significance of SSM, we conducted a similar ablation study by replacing SSM blocks in CodeSSM with DFT (see Figure˜2), creating a variant we call CodeF. Furthermore, we analyze the importance of bidirectional processing via CodeSSM-Uni (Unidirectional Gated SSM), a variant of CodeSSM that removes flip operations from the CodeSSM layer.
@@ 2.2 Pretraining
Dataset. We initially trained CodeSSM and a BERT-like transformer model, which we refer to as BertCoder, on 1.8M text code pairs from the CodeSearchNet (CSN) dataset. The small pretraining dataset allows us to quickly experiment with various model configurations. Both models were trained with a context window of 2048, enabling them to be fine-tuned on benchmarks with lengths shorter than 2048. We refer to the CodeSSM model trained with the CSN dataset as CodeSSM-base. We also train CodeSSM variants such as CodeSSM-pos, CodeSSM-do, CodeF and CodeSSM-Uni (see 2.1) on the same dataset and context length.
Context Length and Length Extrapolation To investigate the length extrapolation capability of SSMs, we pre-train CodeSSM with a context window of 256 length. Since we cannot change the context window for BertCoder once pretrained with a smaller context, we pretrained a RoFormer model Su et al. (2024), which is similar to the BERT model but with Rotatory Positional Embeddings (RoPE), with a context window of length 256. We refer to this model as RoCoder.
The CSN dataset separates the docstring and code, which creates an issue when they are concatenated and subsequently truncated. This can result in one modality (text) being much longer than the other modality (code). We also observed that this discrepancy results in a reduction in performance in tasks involving two sequences (refer to column 3 of Table˜2).
To alleviate this issue, we use a subset of 1.8M samples from the StarCoder dataset (Li et al., 2023) to train models with 256-length context (referred to as CodeSSM-sc-256). This subset contains 300k samples from the six programming languages (PLs) in the CSN dataset  [^2] Go, Java, Javascript, PHP, Python, Ruby. We also train CodeSSM on this dataset with a context length of 2048 (referred to as CodeSSM-sc-2048).
Data Scaling. To study the impact of large datasets on the downstream performance of CodeSSM, we also train it in a large subset of the StarCoder dataset with 4.8M samples (referred to as CodeSSM-sc-L-2048). This subset contains 600k samples from 8 PLs (C and C# along with CSN PLs).
@@ 3 Experiments
In this section, we describe baselines to which we compare CodeSSM and the benchmarks on which we evaluate it.
@@ 3.1 Baselines
We trained two transformer-based models as baselines: a large context (2048) BERT model to compare SSMs’ large context abilities and a small context RoFormer model to compare the length extrapolation abilities of transformers and SSMs. We also train a Zamba model to compare our model with the other SSM variants. The models are trained in the same settings: the same data set, the same objective, the same context length, and batch size.
Additionally, we also mention the performance of some well-performing transformer models in each task to place CodeSSM in the context of SOTA encoder-only transformer models. However, it should be noted that these transformer models are trained on a much larger dataset and with a mixture of multiple training objectives (see Table˜1).
@@ 3.2 Benchmarks
Retrieval tasks. There are two retrieval tasks on which we evaluate our models, NLCodeSearch (AdvTest) Lu et al. (2021) and Stackoverflow Question Answer (SQA). In NLCodeSearch, a description in natural language is provided, and the goal is to find the source code that corresponds to that description. To evaluate a model’s ability to generalize, function names and variable names in the test sets are replaced with generic tokens, for example, the function name with func. This task is difficult because the model has to locate the correct code throughout the test set. In SQA, a StackOverflow question is given, and the goal is to retrieve the highest upvoted answer. This dataset is derived from the original StackOverflow dataset  [^3] https://www.kaggle.com/datasets/stackoverflow/stacksample/data. The average lengths of question and answer are 1.2k and 1.4k, respectively.
Sequence classification tasks. Sequence classification includes two vulnerability detection tasks, Devign Zhou et al. (2019) and DiverseVul Chen et al. (2023a), as well as a complexity prediction task Jeon et al. (2023). Vulnerability detection is a binary classification task in which the goal is to identify whether a given code contains security vulnerabilities. The Devign benchmark Zhou et al. (2019) is a balanced dataset, while DiverseVul consists of only around 5% vulnerable samples. So, we use accuracy for the former and F1 macro for the latter. Complexity prediction Jeon et al. (2023) is a small benchmark with only 3613 training samples to predict the algorithmic complexity of a given Java code among 7 different labels.
Token classification task. In the type-inference task Jesse and Devanbu (2022), the model has to predict the type (such as integer, boolean, etc.) of each token in a given code. The label consists of 50K most frequent (built-in and user-defined) types, while the remaining types have been replaced with the UNK type. During the evaluation, predicting the UNK token is considered incorrect.
@@
For all tasks, we maintain the same context length as specified in the original benchmarks. All models are fine-tuned using a context length of 512 for text-to-code search, clone detection, type inference, and DiverseVul. For the Devign benchmark, the models are fine-tuned with a context length of 400. In SQA and complexity prediction, we use a context length of 1024.
@@
Table 3: Results on SQA benchmark in terms of MRR. The best performance is in bold and other noteworthy results are underlined.
@@ 4 Results
We present the performance of our model on various benchmarks described in Section˜3.2
@@ 4.1 Retrieval Tasks
NLCodeSearch(AdvTest). Column 3 of Table˜2 presents the results of various models evaluated in NLCodeSearch benchmark. Among the models trained on the CSN dataset, CodeSSM-base (with a context size of 2048) outperforms other models, including CodeBERT, which was pretrained with significantly more data. CodeSSM-base with a context length of 256 performs poorly compared to other variants – we explain the reason for this in Section˜2.2. CodeSSM-sc-256 shows remarkable performance and length extrapolation to double the length on which it was trained and is competitive with RoCoder and CodeSSM-sc-2048. However, Zamba – the other SSM based model – failed to train on retrieval task. We attribute the inability to train on retrieval-based tasks to the causal convolution in the Mamba block, which prevents bidirectional understanding.
@@
StackOverflowQA. Table˜3 presents the performance of CodeSSM in conjunction with the baseline models trained by us on the SQA benchmark – BertCoder and RoCoder. SQA allows us to evaluate the models with longer contexts. CodeSSM-sc with a pretraining context length of 256 achieves strong results: it successfully extrapolates to contexts four times longer than those seen during pretraining. Furthermore, CodeSSM-sc outperforms BertCoder (pretrained with a context length of 2048) by 6 points, despite being pretrained with a context length eight times shorter. CodeSSM also surpasses RoCoder, highlighting its superior ability to extrapolate to longer contexts and capture long-range dependencies. Furthermore, we observe that increasing the pretraining context length leads to further improvements in CodeSSM’s performance. Similarly to NLCodeSearch, Zamba did not train on SQA as well.
@@ 4.2 Sequence classification tasks
Devign. Column 3 of Table˜4 shows the results of the various models on the devign benchmark. CodeSSM outperforms BertCoder and RoCoder as well as CodeBERT, GraphCodeBERT, and CodeT5. Interestingly, BertCoder, pretrained with a longer context, also outperforms CodeBERT, showcasing the importance of a longer context for transformers. However, in the case of CodeSSM, the smaller context model performs much better. Notably, RoCoder performs poorly, potentially due to overfitting, as the devign dataset is very small (only 21k training samples). RoPE, used in RoCoder, is also prone to positional overfitting Liu et al. (2024b); Goel et al. (2025). Despite being a unidirectional model, Zamba performs slightly better than RoCoder.
DiverseVul. Column 2 of Table˜4 shows that CodeSSM and its variants outperform BertCoder and RoCoder on the DiverseVul benchmark, which is a larger vulnerability detection dataset compared to Devign. It shows that CodeSSM can perform better than transformers even on a larger dataset. CodeSSM even outperforms CodeBERT, GraphCodeBERT, and CodeT5, which are trained on a much larger dataset and have multiple pretraining objectives (see Table˜1).
Complexity Prediction. Table˜5 shows that CodeSSM-sc-256 outperforms all transformer models, even those trained with a longer context length (BertCoder) and larger data and more pretraining objectives (CodeBERT, GraphCodeBERT, UnixCoder). Additionally, CodeSSM-sc-2048 shows very strong performance as pretraining on longer context length improves the performance further on this task.